OpenStudy (anonymous):

A regtangle lot has a length of 15ft and a area of 300 sg ft. Find the width of the lot?

OpenStudy (anonymous):

A = lw A/l = w

OpenStudy (anonymous):

Width=Area/Length =300/15 =20